when two vectors acting in the same direction are added tge give the resultant of 70 unit when they act in perpendicular direction they give a resultant 50 unit what will be the sum of 2 units if they were acting in the opposite direction


1
Expert's answer
2021-03-08T08:17:26-0500
"a+b=70\\\sqrt{a^2+b^2}=50\\\sqrt{a^2+(70-a)^2}=50\\a=30\\b=40"

Thus,


"b-a=40-30=10"